OpenStudy (anonymous):

A woman invested 4 times as much at 8% as she did at 11%. The total amount of interest she earns in 1 year from both accounts is $430. How much did she invest at each rate

OpenStudy (anonymous):

\[.08\times 4x+.11\times x = 430\] solve for x

OpenStudy (anonymous):

i would clear the decimals by multiplying both sides by 100 to make life easier \[8\times 4x+11x=43000\] \[43x=43000\] \[x=1000\]

OpenStudy (anonymous):

@satellite - have you done Lagrange Multipliers before?